# SuPreME: A Supervised Pre-training Framework for Multimodal ECG Representation Learning

### v1

@@ A.3 Performance Evaluation Metric
We use zero-shot learning and linear probing to evaluate the performance of SuPreME and mainstream eSSL frameworks. The primary metric is Area Under the Receiver Operating Characteristic (AUROC, also referred to as AUC). AUROC is widely used to evaluate the performance of binary classification models. The ROC curve plots the True Positive Rate (TPR) on the vertical axis against the False Positive Rate (FPR) on the horizontal axis. By varying the classifier’s threshold, $\mathbf{TPR}$ and $\mathbf{FPR}$ are calculated and then plotted to form the curve, where $\mathbf{TP}$ refers to True Positive, $\mathbf{FN}$ refers to False Negative, $\mathbf{FP}$ refers to False Positive, and $\mathbf{TN}$ refers to True Negative.:
@@
A.4 Overlap Analysis
Table 6 shows the overlap between entities from the pre-train dataset and cardiac queries from the downstream datasets, filtered using a cosine similarity threshold of 0.95.
Pre-train Dataset Entity Downstream Cardiac Query Similarity Score
@@ A.4 Overlap Analysis
ST-T change st-t changes 0.9968
@@
left ventricle hypertrophy left ventricular hypertrophy 0.9924
@@
supraventricular arrhythmia supraventricular tachycardia 0.9684
Table 6: Overlap between pre-train dataset entities and downstream cardiac queries, filtered with similarity threshold of 0.95, sorted by similarity score.
A.5 Dataset Overview
MIMIC-IV-ECG. MIMIC-IV-ECG [^9] https://physionet.org/content/mimic-iv-ecg/1.0/. is a comprehensive database containing 800,035 diagnostic ECG samples from 161,352 unique patients, with 12-lead recordings in 10 second length and sampled at 500 Hz (Gow et al., 2023). These data have been matched with patient records in the MIMIC-IV clinical database, allowing for the association of waveforms with reports when a cardiologist’s report is available through provided linking information. To enhance the usability of the data, we exclude empty reports as well as reports containing fewer than 3 words, and replace ’NaN’ and ’Inf’ values in the ECG records with the average of 6 neighboring points. Ultimately, the dataset used for clinical entity extraction tasks includes 771,500 samples, each comprising 18 machine-generated ECG reports based on rules and the corresponding ECG data. After clinical NER and dedplication on the 18 ECG reports of each sample, the dataset holds 295 labels of professional medical terminologies.
PTB-XL. PTB-XL [^10] https://physionet.org/content/ptb-xl/1.0.3/. is a large open-source ECG dataset, comprising 21,799 clinical ECG records from 18,869 patients, with each lead sampled at a rate of 500 Hz and a duration of 10 seconds (Wagner et al., 2020). A total of 71 different ECG reports are SCP-ECG compliant, covering diagnostic, form and rhythm reports. PTB-XL also provides a recommended train-test split and includes multi-level ECG annotations, covering Superclass (5 categories), Subclass (23 categories), Form (19 categories), and Rhythm (12 categories). Notably the 4 subsets have different sample sizes.
CPSC-2018. The CPSC-2018111111http://2018.icbeb.org/Challenge.html. dataset originates from the China Physiological Signal Challenge (CPSC) 2018, including 6,877 records from 9,458 patients, with durations ranging from 6 to 60 seconds (Liu et al., 2018). The standard 12-lead ECG data is sampled at a rate of 500 Hz, collected from 11 hospitals and categorized into 9 different labels: 1 normal type and 8 abnormal types.
Chapman-Shaoxing-Ningbo (CSN). The CSN [^12] https://physionet.org/content/ecg-arrhythmia/1.0.0/. 12-lead ECG dataset is created with the support of Chapman University, Shaoxing People’s Hospital and Ningbo First Hospital, which includes 12-lead ECGs from 45,152 patients, with a sampling rate of 500 Hz and a duration of 10 seconds (Zheng et al., 2020, 2022). It contains expert annotated features that cover variety of common heart rhythms and other cardiovascular conditions. We exclude ECG records with "unknown" annotations and get 23,026 ECG records with 38 different labels.
A.6 Downstream Data Split
@@
A.7 Downstream Experiment Configuration
@@
Type Cosine Anealing Cosine Anealing Cosine Anealing Cosine Anealing Cosine Anealing Cosine Anealing
@@
A.8 Performance on Linear probing.
Table 9 shows the linear probing AUC performance of SuPreME’s and other eSSLs’ ECG encoders on specific six downstream datasets.
From Scratch
ECG Only
Multimodal Learning
MERL (ResNet) 82.39 86.27 88.67 64.90 80.56 84.72 58.26 72.43 79.65 53.33 82.88 88.34 70.33 85.32 90.57 66.60 82.74 87.95
MERL (ViT) 78.64 83.90 85.27 61.41 77.55 82.98 56.32 69.11 77.66 52.16 78.07 81.83 69.25 82.82 89.44 63.66 78.67 84.87

### v2

@@ A.5.3 Evaluation Metrics
We use zero-shot learning and linear probing to evaluate the performance of our framework and mainstream eSSL frameworks. The primary metric is Area Under the Receiver Operating Characteristic (AUROC, also referred to as AUC). AUROC is widely used to evaluate the performance of binary classification models. The ROC curve plots the True Positive Rate (TPR) on the vertical axis against the False Positive Rate (FPR) on the horizontal axis. By varying the classifier’s threshold, $\mathbf{TPR}$ and $\mathbf{FPR}$ are calculated and then plotted to form the curve, where $\mathbf{TP}$ refers to True Positive, $\mathbf{FN}$ refers to False Negative, $\mathbf{FP}$ refers to False Positive, and $\mathbf{TN}$ refers to True Negative.:
@@
A.6 Statistics and Overlap Analysis
A.6.1 Statistics of Extracted MIMIC-IV-ECG Entities
We extract over 3.4 million clinical entities from free-text ECG reports in the MIMIC-IV-ECG dataset using an instruction-tuned LLM. At the term level, this results in 1,168 unique raw entities (Table 13). Among these, 93.75% remain after filtering out uncertain or ambiguous expressions. To resolve redundancy and lexical variation, we apply embedding-based clustering using MedCPT representations, reducing the vocabulary to 341 cluster representatives. Further manual verification and mapping to UMLS/SNOMED CT terminologies yield a final set of 295 standardized cardiac entities used as global queries during supervised pre-training.
Entity Type Count Proportion
Raw extracted entities 3,419,064 100% (sample-level)
Unique raw extracted entities 1,168 100% (term-level)
Terms after uncertainty filtering 1,095 93.75% (vs. 1,168)
Entity cluster representatives (post-deduplication) 341 29.20% (vs. 1,168)
Final unique standardized entities (post-mapping) 295 25.26% (vs. 1,168)
Table 13: Statistics of unique cardiac Entities: Extraction, Filtering, Deduplication, and Mapping
Table 14 provides additional statistics on the clustering process. The average cluster contains 3.39 entities, with some clusters merging up to 29 semantically similar terms. In total, 86.51% of clusters are successfully mapped to standardized terms. The distribution of standardized entity frequencies is illustrated in Figure 10. The left panel shows a log-scaled histogram of the most common cardiac terms, with "normal", "abnormal", and "myocardial infarction" being the most frequent. The right panel presents a word cloud that qualitatively reflects term prevalence and semantic variety. Together, these visualizations confirm that while a few diagnostic terms dominate the corpus, a long tail of clinically significant but less frequent entities is preserved, supporting robust coverage in downstream classification.
Clustering Metric Value
Number of clusters formed 341
Average number of entities per cluster 3.39
Maximum / Minimum cluster size 29 / 1
Proportion of clusters mapped to standard terms 86.51%
Table 14: Clustering statistics of extracted cardiac entities on MedCPT embeddings
Figure 10: Frequency distribution of standardized ECG entities after deduplication and mapping in MIMIC-IV-ECG.
A.6.2 Dataset Overlap Analysis
We analyze the cardiac query overlap between the pre-training dataset and six downstream datasets specified in Section 4.1, as well as among the downstream datasets themselves, as illustrated in Figure 11. Specifically, we embed all entities from the pre-training dataset and cardiac queries from the downstream datasets, compute their cosine similarity, and apply a threshold of 0.95 verified by cardiologists with 10+ years of experience as well to filter overlapping queries.
Figure 11: Overlap between ECG datasets, with left panel showing pairwise overlap counts between downstream datasets, and right panel showing the distribution of overlapping and non-overlapping classes between each downstream dataset and the pre-training dataset.
The heatmap on the left shows that pairwise overlaps among downstream datasets are generally limited, reflecting the diversity of cardiac query prompts. The bar chart on the right reveals that 57 cardiac queries overlap between the pre-training dataset and the downstream datasets. Despite the pre-training dataset shares some similar queries, a substantial portion of queries remains unique to the downstream datasets, allowing the pre-training process to establish robust general-purpose representations while leaving room for downstream-specific adaptation.
Table 15 shows the overlap between entities from the pre-training dataset and cardiac queries from the downstream datasets, filtered using a cosine similarity threshold of 0.95.
Pre-training Dataset Entity Downstream Cardiac Query Similarity Score
@@ A.6.2 Dataset Overlap Analysis
left ventricle hypertrophy left ventricular hypertrophy 0.9924
@@
Q wave present q wave 0.9903
@@
Table 15: Overlap between pre-training dataset entities and downstream cardiac queries, filtered with similarity threshold of 0.95, sorted by similarity score.
A.7 Downstream Task Configuration
A.7.1 Downstream Data Split
@@
A.7.2 Downstream Experiment Configuration
@@
Type Cosine Annealing Cosine Annealing Cosine Annealing Cosine Annealing Cosine Annealing Cosine Annealing
@@
A.8 Performance of Non-overlapped Cardiac Conditions
While evaluating performance exclusively on non-overlapping (i.e., unseen) downstream classes is not a standard practice in existing ECG literature, including MERL and other multimodal or self-supervised frameworks, we acknowledge its value in assessing true generalization. To address this, we conduct an additional analysis where we evaluate zero-shot AUC only on downstream classes that do not appear in the pre-training dataset.
Table 18 presents the comparison between AUC scores on all downstream classes versus only non-overlapping ones. As expected, performance on unseen classes is moderately lower, yet remains strong across datasets, confirming our framework’s ability to generalize beyond pre-trained diagnostic categories. This analysis complements our main results and provides deeper insights into model robustness.
Setting PTB-XL-Super PTB-XL-Sub PTB-XL-Form PTB-XL-Rhythm CPSC-2018 CSN Overall
Non-overlapping Classes 75.97 69.30 61.36 83.83 75.73 73.24
All Classes 78.20 77.52 60.67 86.79 79.83 80.17 77.20
Table 18: Zero-shot AUC on downstream datasets using only non-overlapping (unseen) classes vs. using all classes.
@@ A.9 Performance on Specific Cardiac Conditions
A.10 Further Discussion
A.10.1 Offline Use of Large-sized LLM for Clinical NER
While we employ LLaMA3.3-70B-Instruct for clinical entity extraction, this step is performed offline only once during dataset construction and is not part of the SuPreME model’s pre-training & inference pipeline. The motivation for using a larger model is to ensure high annotation quality for the pre-training dataset. Once entities are extracted and mapped, they form a standardized query list used throughout training and evaluation. Therefore, clinical deployments do not require access to large LLMs, and the SuPreME model itself remains lightweight during inference.
A.10.2 Computation Cost and Practical Deployment
Data Processing (Offline NER). To obtain high-quality supervision labels, we extract and normalize diagnostic entities from MIMIC-IV-ECG reports using LLaMA3.3–70B-Instruct with structured prompts. This step is performed only once as described above before pre-training SuPreME model. The output is a cleaned, deduplicated dataset of standardized diagnostic labels, which serves as supervision for training. The annotation process takes approximately 6 hours on 8 NVIDIA A100-SMX4-80GB GPUs, and the tested minimum reproducing resources are 4 NVIDIA A100-PCIE-40GB GPUs without parallelized LLM inference. The resulting standardized dataset is reused across training and downstream evaluation.
SuPreME Pre-training. SuPreME itself consists only of a ViT-based ECG encoder, query-based supervision, and a lightweight Cardiac Fusion Network (CFN). Training is efficient, around 1.5 hours on 4 NVIDIA A100-PCIE-40GB GPUs achieving best AUC performance (16 epochs), and does not require contrastive sampling or further fine-tuning in deployment.
Deployment and Inference Once pre-trained, SuPreME supports zero-shot ECG classification via a set of concise cardiac query prompts. Inference only involves a forward pass through the ECG encoder and CFN, taking milliseconds per ECG sample. No LLMs or textual reports are needed at test time, making SuPreME highly practical for deployment in real-world clinical settings. We empirically verify that inference can be efficiently performed on a single NVIDIA A5000-PCIE-24GB GPU or NVIDIA RTX4090-PCIE-24GB GPU.
A.10.3 Similarity Threshold Determination
The similarity thresholds in our entity deduplication and mapping pipeline were determined in consultation with experienced cardiologists (over 10 years of clinical practice), based on joint analysis of the results under various threshold settings in each phase.
Through this process, we observed that setting the thresholds too high (e.g., above 0.9 in entity mapping) would exclude valid clinical variants due to minor wording differences, while setting them too low (e.g., below 0.7 in entity mapping) could introduce semantic ambiguity by incorrectly matching unrelated conditions (Table 19).
Standard Terminology Report Entity Similarity Score
left ventricle hypertrophy Ventricular fibrillation 0.6400
non-specific ST changes ST elevation 0.6343
inferior myocardial infarction anterior wall abnormality 0.5717
Table 19: Incorrect matching examples between standard terminology and report entities.
"left ventricle hypertrophy" and "ventricular fibrillation" shows a similarity score of 0.64, but are entirely unrelated - one refers to structural enlargement of the left ventricle, while the other refers to a life-threatening arrhythmia. The selected thresholds reflect a balance between preserving clinically meaningful variants and minimizing noise.
A.10.4 Comparing SuPreME with MERL
To assess the relative effectiveness of our supervised multimodal framework, we compare SuPreME against MERL (Liu et al., 2024), a recent multimodal contrastive learning baseline that utilizes clinical reports and enhanced prompt engineering (Figure 18). While MERL employs contrastive objectives and handcrafted prompts, SuPreME leverages fine-grained diagnostic supervision through LLM-extracted entities and multimodal fusion via the Cardiac Fusion Network (CFN).
Figure 18: Per-dataset AUC difference between SuPreME and MERL.
Framework PTB-XL-Superclass PTB-XL-Subclass PTB-XL-Form PTB-XL-Rhythm CPSC-2018 CSN Avg
MERL 73.89 74.32 61.54 79.89 76.01 75.61 73.54
SuPreME 78.20 77.52 60.67 86.79 79.83 80.17 77.20
Table 20: Zero-shot AUC comparison between SuPreME and MERL.
As shown in Table 20, SuPreME achieves consistently higher AUCs across all but one dataset. On average, SuPreME improves zero-shot performance by 3.66% absolute. Statistical testing confirms this improvement is significant: a paired $t$-test across the six datasets yields $t=3.51$, $p=0.0171$.
Framework Zero-shot AUC (%)
MERL $73.54\pm 2.30$
SuPreME (Ours) $\mathbf{77.20\pm 0.21}$
Table 21: Average zero-shot AUC and standard deviation across six datasets.
Moreover, SuPreME demonstrates significantly lower performance variance across datasets. While MERL exhibits a standard deviation of 2.30, SuPreME achieves a much smaller deviation of 0.21 (Table 21), indicating greater robustness and stability across diverse cardiac classification tasks. These results collectively support the effectiveness of our proposed supervised pre-training framework and its entity-level modality fusion strategy, even when compared to a strong multimodal baseline.
A.10.5 On the Effectiveness of CFN with Different Backbones
To address concerns regarding the effectiveness of the Cardiac Fusion Network (CFN), particularly its relatively lower performance when paired with a ResNet backbone (cf. Table 2), we conduct statistical analyses to better understand the interaction between backbone architecture and the CFN module.
$\Delta$ AUC Comparison. We compare the AUC improvement brought by CFN over linear classification for both ViT and ResNet backbones across six downstream datasets. As shown in Figure 19, CFN brings consistent performance gains when combined with ViT, with average improvement of +5.97 AUC. In contrast, CFN shows little to negative improvement with ResNet, indicating that the quality of the underlying feature representations plays a critical role in effective cross-modal fusion.
Figure 19: CFN vs. linear classification ($\Delta$ AUC) for ViT ECG backbone.
Statistical Significance. To verify this trend, we perform a paired t-test and Wilcoxon signed-rank test on the $\Delta$ AUC values. Both tests confirm that CFN yields significantly greater improvements on ViT than ResNet:
  • Paired t-test: $t=4.99$, $p=0.0021$
  • Wilcoxon test: $W=21.0$, $p=0.0156$
These results provide strong statistical evidence that ViT synergizes better with CFN compared to ResNet, likely due to ViT’s superior capacity in capturing global temporal dependencies in ECG signals.
CFN is designed to align high-level ECG features with cardiac queries via cross-attention. However, ResNet provides only local, convolutional features with limited contextual depth, especially compared to ViT’s global receptive field. As a result, the decoder lacks sufficient global representations to effectively condition on query semantics. This bottleneck explains the performance drop observed in ResNet + CFN.
Two-Way ANOVA. We further conduct a two-way ANOVA with Backbone (ResNet vs. ViT) and Module (Linear vs. CFN) as factors. As shown in Table 22, the interaction term is statistically significant ($F=6.60$, $p=0.018$), confirming that the effect of CFN depends on the choice of backbone. Notably, neither factor alone is significant, suggesting that their combination determines performance.
Source Sum of Squares df F-value p-value
Backbone 167.06 1 3.79 0.066
Module 5.57 1 0.13 0.726
Backbone $\times$ Module 290.65 1 6.60 0.018
Residual 881.22 20 - -
Table 22: Two-way ANOVA results on AUC with backbone and module as factors.
Takeaway. These findings reinforce CFN’s role as a powerful fusion mechanism when paired with a backbone (like ViT) that produces expressive feature sequences. The drop in performance with ResNet may stem from its less structured output, which lacks the sequential token-style organization needed for effective query-based attention. Thus, the CFN is not inherently ineffective, but its utility hinges on a compatible encoder design.
A.10.6 Domain Scope and Generalization Potential
While our framework is evaluated on 12-lead ECG data, we believe that this modality represents a highly impactful and widely applicable domain in clinical practice. ECG is routinely used across diverse medical contexts, including emergency rooms, intensive care units (ICUs), outpatient cardiology clinics, and even home-based healthcare monitoring, due to its low cost, non-invasiveness, and real-time ability to reflect cardiac electrical activity. As such, improving automated ECG interpretation has direct clinical relevance across resource settings and specialties.
Moreover, although this work focuses on ECG, the core methodology of SuPreME, namely multimodal learning between biomedical signals and clinically meaningful queries, can be generalized to other physiological signal domains such as EEG (electroencephalogram) or PPG (photoplethysmography). These modalities are similarly structured (multi-channel, time-series signals) and increasingly available in clinical and wearable settings. However, to the best of our knowledge, there is currently a lack of large-scale, publicly accessible datasets that pair these signals with detailed, free-text clinical reports suitable for training our entity extraction module.
We hope our work can inspire future efforts toward building such paired datasets for other biomedical signals, enabling the broader application of query-based multimodal learning frameworks beyond ECG.